The seats of a jet airliner are arranged in 55 rows(numbered 1-55) of 10 seats (lettered A to K, leaving out I).
In each row, seats C, D, G and H are on aisles, and A and K are window seats. Smoking is permitted in row
45 to 55 inclusive.
If a passenger is assigned a seat at random, what is the probability of being allocated
a) An aisle seat?

b) A seat in the smoking section?

c) A window seat in the non-smoking section?

c) A window seat in row 1?
1
Expert's answer
2021-09-12T23:51:40-0400

a) There are 4 aisle seats in each row of 10 seats. "P=\frac{4}{10}=0.4."


b) There are 11*10 seats in smoking section. "P=\frac{11*10}{55*10}=\frac{11}{55}=0.2."


c) There are (55-11)*2 window seats in non-smoking section. "P=\frac{44*2}{55*10}=\frac{88}{550}=0.16."


d) There are 2 window seats in row 1. "P=\frac{2}{550}=\frac{1}{275}\approx0.0036."